A vertical pipe opened at both ends is immersed into a liquid. The length of the air column in the pipe is l. A tuning fork of frequency 342 Hz is held over the open end of the pipe while it is gradually being raised. The 1st and 3rd harmonics occur in air at the length of 235 mm and 744 mm respectively. Calculate (i)The wavelength of the sound (ii) The speed of the sound in air (iii) The end correction
1
Expert's answer
2020-11-17T13:04:43-0500

i)


"L=\frac{\lambda}{4}\\0.235=\frac{\lambda}{4}\\\lambda=0.94\ m"

ii)


"v=\lambda f=(342)(0.94)=321\frac{m}{s}"

"x=vT=v\frac{2n-1}{f}=321\frac{2(2)-1}{342}=2.82\ m"